if the number of turns becomes double but the length remain same then magnetic field in the solenoid becomes???


1
Expert's answer
2021-05-24T14:09:36-0400

Gives 

solenoid


Case(1)

length

"l_1=l"

Turn


"N_1=n"

Magnetic field

B=B1

case(2)

length

"l_2=l"

Turn


"N_2=2n"

Magnetic field

B=B2


We know that

Magnetic field in solenoid

"B=\mu_0\frac{N}{l}i"

n=turn per unit length

we can written as 

"\frac{B_1}{B_2}=\frac{N_1}{N_2}\frac{l_2}{l_1} \rightarrow (1)"

Put value

"\frac{B_1}{B_2}=\frac{n}{2n}\frac{l}{l}"

"\frac{B_1}{B_2}=\frac{1}{2}"

"B_2=2B_1"

Final Magnetic field is equal to double of initial magnetic field